OpenStudy (anonymous):

Samantha deposits $5,000.00 in account that earns 5.5% simple interest. About how many years will it take for her account balance to earn $3,000.00 in interest?

OpenStudy (anonymous):

will give a metal

OpenStudy (chrperez28):

SI = PxTXR/100 3000=5000xTx5.5/100 3/5x100/5.5=T 10 .9 years

OpenStudy (anonymous):

the possible answers are About 2 years About 5 years About 11 years About 20 years

OpenStudy (chrperez28):

About 11 years

OpenStudy (anonymous):

thank you @chrperez28

OpenStudy (kash_thesmartguy):

\[Interest=Principal*Rate*Time\]\[3000=5000*0.055*x\]\[3000=275x\]\[\frac{ 3000 }{ 275 }=\frac{ 275x }{ 275 }\]\[x=10.9\]About 11 years.

OpenStudy (anonymous):

thank you @Kash_TheSmartGuy

OpenStudy (kash_thesmartguy):

no prob :)